# Lower Bounds for Differential Privacy Under Continual Observation and Online Threshold Queries

### v1

@@ B.2 JDP of Algorithm 2
It follows from Claim 3 that for any $j\geq 1$, $\Pr[X=j+1]\leq e^{0.5}\Pr[X=j]$. For $\varepsilon<0.6$ we have $\varepsilon e^{0.5}<1$.
We specify the following predicate $F$ for excluded runs: $F$ holds when $K_{1}\in\{1,L\}$ or when $K_{2}\in\{1,L\}$ or when the output (on compressed input) has the general form where for some $i$, $**\boldsymbol{y}_{-i}=\bot^{*}\top^{*}$.
Claim 5 (Specified Excluded Runs).
For any input with at least $2L$ steps (where $L$ is as in Claim 3) with $\Delta_{t}=1$, the probability of predicate $F$ is $O(\delta)$.
Proof B.9.
The events $K_{1},K_{2}\in\{1,L\}$ have total probability at most $4\delta$. The probability of an output of the form $**\boldsymbol{y}_{-i}=\bot^{*}\top^{*}$ is at most that that during Phase II, all the steps with $\Delta_{t}=1$ have $y_{t}=\top$ with the exceptions: When $i$ is in the $\bot$ sequence, then only starting at count $C_{2}=2$. Or when $i$ is in the $\top$ sequence then step $t=i$ is not required to have $y_{t}=\top$. This probability is for $k\in\{1,2\}$, at most $\sum_{j\geq k}\Pr[K_{2}=i]\varepsilon^{j-1}=O(\delta)$, using Claim 4.
We fix an output sequence $\boldsymbol{a}_{-i^{*}}$ in all positions but $i^{*}$. In order to establish (6) it suffices to establish that
$\Pr[\boldsymbol{y}_{-i^{*}}=\boldsymbol{a}_{-i^{*}}]\approx_{O(\varepsilon)\mid F}\Pr[\boldsymbol{y}^{\prime}_{-i^{*}}=\boldsymbol{a}_{-i^{*}}].$ (7)
When $\boldsymbol{a}_{-i^{*}}$ is such that $\Pr[\boldsymbol{y}_{-i^{*}}=\boldsymbol{a}_{-i^{*}}]=0$ and $\Pr[\boldsymbol{y}^{\prime}_{-i^{*}}=\boldsymbol{a}_{-i^{*}}]=0$ then (7) clearly holds. We therefore assume going forward that this is not the case.
Let
$w_{1}$ :=$\arg\min_{t\not=i^{*}}a_{t}=\top$
$w_{2}$ :=$\arg\max_{t\not=i^{*}}a_{t}=\bot$
$A$ $:=\sum_{t\in[w_{1}:w_{2}]\setminus\{i_{*}\}}\boldsymbol{1}(a_{t}=\top)$
respectively, be the position of the first $\top$” in $\boldsymbol{a}_{-i^{*}}$, the position of the last $\bot$” in $\boldsymbol{a}_{-i^{*}}$, and the number of $\top$” occurrences in the positions $\{w_{1},\ldots,w_{2}\}$.
Note that there can be at most $2L$$\top$” occurrences before $w_{2}$. It holds that $w_{1}\leq w_{2}-2$ (and equality is possible only when $i^{*}=w_{1}+1$). It holds that $t_{1}\leq w_{1}$, that is, Phase I must end before $t=w_{1}$. It also holds that $t_{2}>w_{2}$. That is, Phase III can start only when $t>w_{2}$. In particular for $w_{1}\leq t\leq w_{2}$ we are at Phase II. The case $A=0$ corresponds to $w_{2}\leq w_{1}$.
It suffices to establish (7) conditioned on fixing the randomness $\boldsymbol{r}_{-i^{*}}$ of the activation bits in all time steps $t<w_{1}$ or $t\not=i^{*}$. Let $R:=\sum_{t\in[w_{1}-1]\setminus\{i^{*}\}}r_{t}$ be the number of activated steps before $w_{1}$. Note that throughout Phase II, and in particular for $w_{1}\leq t\leq w_{2}$ ($t\not=i^{*}$), $r_{t}$ is determined by $a_{t}$, that is, $r_{t}=1\iff a_{t}=\top$. The probability of that is $Q_{A}=\varepsilon^{A}$ The remaining randomness in $\boldsymbol{y}_{-i^{*}}$ we need to consider when computing the probability of output $\boldsymbol{a}_{-i^{*}}$ is due to $r_{i^{*}}$, $r_{w_{2}+1:T}$, $K_{1}$, and $K_{2}$.
Case $A=0$ ($w_{2}\leq w_{1}$)
In this case $\boldsymbol{a}_{-i^{*}}$ has the form $\bot^{*}\top^{*}$ and $F\equiv 1$ conditioned on this output. Therefore (7) holds.
Case $A=1$ ($w_{2}>w_{1}$)
With input $\boldsymbol{\Delta^{\prime}}$ we have $\Delta^{\prime}_{i^{*}}=0$. We must have $K_{1}=R$, $C_{2}(w_{2})=A$, and $K_{2}>A$. $K_{2}=A+j$ is possible only if we get samples of $r_{t}=1$ in the first $j$ steps that are not $i^{*}$ starting from $w_{2}+1$. Hence,
$\Pr[\boldsymbol{y}^{\prime}_{-i^{*}}=\boldsymbol{a}_{-i^{*}}]$ $=\Pr[K_{1}=R]\cdot\varepsilon^{A}\cdot\sum_{j=1}^{L-A}\varepsilon^{j}\Pr[K_{2}=A+j]\ .$ (8)
With input $\boldsymbol{\Delta}$, we consider the following cases according to where $i^{*}$ is in the sequence.
  • [case $i^{*}>w_{2}$] It is the same calculation as (8) except that $i^{*}$ contributes in the last factor, replacing the contribution of the next time step and so forth. There is no difference if the sequence is long enough $T>w_{2}+L-A$ as assumed. Therefore (7) holds with equality.
  • [case $w_{1}+1\leq i^{*}\leq w_{2}-1$] We know that step $i^{*}$ is in Phase II of the algorithm and the first $\top$ was returned after exactly $R$ inputs with $\Delta_{t}=1$. Hence, this is only possible with $K_{1}=R$. We decompose the probability by conditioning it on the value of the activation $r_{i^{*}}$:
    $\Pr[\boldsymbol{y}_{-i^{*}}=\boldsymbol{a}_{-i^{*}}]$ $=(1-\varepsilon)\cdot\Pr[\boldsymbol{y}_{-i^{*}}=\boldsymbol{a}_{-i^{*}}\mid r_{i^{*}}=0]+\varepsilon\cdot\Pr[\boldsymbol{y}_{-i^{*}}=\boldsymbol{a}_{-i^{*}}\mid r_{i^{*}}=1]$ (9)
    We will establish that
    $\Pr[\boldsymbol{y}_{-i^{*}}=\boldsymbol{a}_{-i^{*}}\mid r_{i^{*}}=0]$ $=\Pr[\boldsymbol{y}^{\prime}_{-i^{*}}=\boldsymbol{a}_{-i^{*}}]$ (10)
    $\Pr[\boldsymbol{y}_{-i^{*}}=\boldsymbol{a}_{-i^{*}}\mid r_{i^{*}}=1]$ $\approx_{O(1)\mid F}\Pr[\boldsymbol{y}^{\prime}_{-i^{*}}=\boldsymbol{a}_{-i^{*}}]$ (11)
    Combining (10) and (11) in (9) establishes the claim (7) for this case.
    When $r_{i^{*}}=0$, $C_{2}$ is not incremented in step $i^{*}$. The calculation is the same as in (8), which establishes (10).
    When $r_{i^{*}}=1$ then $C_{2}$ is incremented at step $i^{*}$. The transition to Phase III still happens at or after step $w_{2}-1$ but at that point we have $C(w_{2}-1)=A+1$. Therefore,
    $\Pr[\boldsymbol{y}_{-i^{*}}=\boldsymbol{a}_{-i^{*}}\mid r_{i^{*}}=1]$ $=\Pr[K_{1}=R]\cdot\varepsilon^{A}\cdot\sum_{j=1}^{L-A}\varepsilon^{j}\Pr[K_{2}=A+j+1]\ .$ (12)
    When $A=L-1$, we exclude the $F$ event of $K_{2}=L$. When $A<L-1$ we apply Claim 4 to obtain that
    $\sum_{j=1}^{L-A}\varepsilon^{j}\Pr[K_{2}=A+j]$ $=O(\varepsilon\Pr[K_{2}=A+1])$
    $\sum_{j=1}^{L-A}\varepsilon^{j}\Pr[K_{2}=A+j+1]$ $=O(\varepsilon\Pr[K_{2}=A+2])\ .$
    Since $\Pr[K_{2}=A+1]\approx_{O(1)}\Pr[K_{2}=A+2]$ (from Claim 3) we obtain that
    $\sum_{j=1}^{L-A}\varepsilon^{j}\Pr[K_{2}=A+j+1]\approx_{O(1)\mid F}\sum_{j=1}^{L-A}\varepsilon^{j}\Pr[K_{2}=A+j]$
    and therefore this establishes (11).
  • [case $i^{*}<w_{1}$] We follow the same method as the prior step and establish that (10) and (11) hold. When $r_{i^{*}}=0$, the calculation is the same as with $\boldsymbol{\Delta}^{\prime}$, as no counter is incremented, and this establishes (10). When $r_{i^{*}}=1$, we consider the following subcases:
    • (i)
[…]
      • (i)A
[…]
      • (i)B
        A transition to Phase III did not occur at $i^{*}$ and hence we have $C_{2}=2$ at step $w_{1}$. We obtain the exact same expression for the conditional probability as (12) and therefore (11) holds for this case as well and the argument proceeds the same way.
    • (ii)
      $i^{*}$ contributed towards the Phase I count and $w_{1}$ is the first count of Phase II. Note that this is possible only when $R<L$. In this case, $K_{1}=1+R$ but otherwise the sequences are “synched” with $C_{2}=1$. We obtain
      $\Pr[\boldsymbol{y}_{-i^{*}}=\boldsymbol{a}_{-i^{*}}\mid r_{i^{*}}=1]$ $=\Pr[K_{1}=R+1]\cdot\varepsilon^{A}\cdot\sum_{j=1}^{L-A}\varepsilon^{j}\Pr[K_{2}=A+j]\ .$ (13)
      Claim (11) for this case follows using that (Claim 3) for $1\leq R<L$, $\Pr[K_{1}=R]\approx_{O(1)}\Pr[K_{1}=R+1]$.
@@ Appendix C Private Online Prediction for Point Functions
    Instantiate the Delayed-Mirror (Remark B.5) variant of Algorithm JDP-Mirror with the privacy parameter $\varepsilon$ and with the threshold $K=20\varepsilon k$ for $k=\Theta(\frac{1}{\varepsilon}\log\frac{1}{\delta})$.
@@
        If ${\rm Count}\geq k=\Theta(\frac{1}{\varepsilon}\log\frac{1}{\delta})$ then
@@
          Run an $(\varepsilon,\delta)$-DP algorithm for sparse histograms to identify a point $x^{*}\in\mathcal{X}$ that appears many times in the multiset $\{x_{j}:j\in[t]\text{ and }y_{j}=1\}$. If the algorithm failed to identify such a point then we assume that $x^{*}=\star$ for some special symbol $\star\notin\mathcal{X}$. See, e.g., Bun et al. (2016); Cohen et al. (2021).
@@
Let $\mathcal{X}$ be a domain. The class of point functions over $\mathcal{X}$ admits an $(\varepsilon,\delta)$-DP online predictor with mistake bound $O(\frac{1}{\varepsilon}\log\frac{1}{\delta})$.
@@
The proof is via the construction given in Algorithm 3, which we denote here as algorithm $\mathcal{A}$. The utility analysis of this algorithm is straightforward: Once $\approx\frac{1}{\varepsilon}\log\frac{1}{\delta}$ positive labels have been obtained, we run the private histogram algorithm in order to identify this positively labeled point and the algorithm does not make anymore mistakes from that point forward.
For the privacy analysis, fix two neighboring input streams that differ on their $i$th labeled example: $S_{0}=((x^{0}_{1},y^{0}_{1}),\dots,(x^{0}_{T},y^{0}_{T}))$ and $S_{1}=((x^{1}_{1},y^{1}_{1}),\dots,(x^{1}_{T},y^{1}_{T}))$. Let $b\in\{0,1\}$ and consider the execution of $\mathcal{A}$ on $S_{b}$. We now claim that the outcome distribution of $\mathcal{A}(S_{b})$, excluding the $i$th output, can be simulated from the outcomes of three private computations on $b$. Specifically, we consider a simulator Sim that interacts with a “helping algorithm”, Helper, who knows the value of the bit $b$ and answers DP queries w.r.t. $b$ as issued by Sim. The simulator Sim first asks Helper to execute the algorithm for histograms on the dataset containing the first $k$ positively labeled points in $S_{b}$ to obtain $x^{*}$. This satisfies $(\varepsilon,\delta)$-DP. The simulator then asks Helper for the outcome vector of executing JDP-Mirror on the bit-stream $\left(I(x^{b}_{j},x^{*})\right)_{j=1}^{T}$, where $I(a,b)=1$ iff $a=b$. The Helper responds with the vector $\vec{b}\in\{\bot,\top,\emptyset\}^{T}$ (where the $i$th coordinate is $\emptyset$, denoting the fact that these coordinate is not given to the simulator). Next, similarly to Step 3(c)ii of the algorithm, let $\hat{L}$ denote the indices of the first $10k$ occurrences of $x^{*}$ in $S_{b}$. The simulator asks Helper for the value of
@@
Now, whenever these two events hold, then the outcome of the simulator is identical to the outcome of $\mathcal{A}(S_{b})$. More specifically, if the algorithm sets ${\rm Flag}=1$ in Step 3iiiC, then the outcome of the simulator is syntactically identical to the output of the algorithm. Otherwise (if we set ${\rm Flag}=3$ in Step 3iiiC), then there must be at least $9k$ occurrences of $(x^{*},0)$ in $S_{b}$ before the $k$th occurrence of $(x^{*},1)$. Hence, in this case we have that $\widehat{{\rm Fake}}\geq\frac{1}{\varepsilon}\log\frac{1}{\delta}$. So both the algorithm and the simulator output the all zero vector.

### v2

@@ Appendix B JDP Algorithm for the Mirror Problem
Fix $i^{*}\in[T]$ and fix two neighboring input sequences $\vec{\Delta},\vec{\Delta}^{\prime}\in\{0,1\}^{T}$ that differ in position $i^{*}$. Observe that coordinates $t$ s.t. $\Delta_{t}=0$ (and hence when $t\not=i^{*}$, $\Delta^{\prime}_{t}=0$) have no effect on either of the executions. Hence, we may ignore these coordinates. Therefore, without loss of generality, we may assume that $\vec{\Delta}\equiv\boldsymbol{1}$, and that $\vec{\Delta}^{\prime}$ is the vector of all 1 except for the $i^{*}$th coordinate for which $\Delta^{\prime}_{i^{*}}=0$.
Consider now for each $t\in[T]\setminus\{i^{*}\}$ the random variables $Y_{t}$ that is the output at time $t$ when the input is $\vec{\Delta}$ and the random variable $Y^{\prime}_{t}$ that is the output at time $t$ when the input is $\vec{\Delta}^{\prime}$. Note that for a fixed input, the random variables of different time steps are independent. Denote $P_{t}=\Pr[Y_{t}=\top]$ and $P^{\prime}_{t}=\Pr[Y^{\prime}_{t}=\top]$. We will show that there can be at most $2L$ time steps $t\in[T]$ where $P_{t}\not=P^{\prime}_{t}$ and that out of these steps, there can be at most two where $|P_{t}-P^{\prime}_{t}|\leq\delta^{\prime}$ (and thus $Y_{t}\approx_{0,\delta^{\prime}}Y^{\prime}_{t}$) and at most $22L-$ 2 where $Y_{t}\approx_{\varepsilon^{\prime}}Y^{\prime}_{t}$. If we apply advanced composition over time steps $t\in[T]\setminus\{i^{*}\}$ we obtain that $\boldsymbol{Y}_{-i^{*}}\approx_{\varepsilon,\delta}\boldsymbol{Y}_{-i^{*}}$ where
$\varepsilon$ $=\frac{1}{2}(2L-2)(\varepsilon^{\prime})^{2}+\varepsilon^{\prime}\sqrt{2(2L-2)\log(1/\delta^{\prime\prime})}=O(\sqrt{\varepsilon^{\prime}}\log(1/\delta^{\prime\prime}))$
$\delta$ $=2\delta^{\prime}+\delta^{\prime\prime}\ .$
Therefore if we wish for final privacy parameters $\varepsilon,\delta$ we need to choose $\varepsilon^{\prime}=O((\varepsilon)^{2}\log^{-2}(1/\delta))$ and $\delta^{\prime}=O(\delta)$.
We can tighten the privacy analysis using the Target Charging Technique (Cohen and Lyu, 2023). We note that the $22T-$ 2 steps that can differ between the dataset have $\top$ probabilities $\pi_{\varepsilon^{\prime},\delta^{\prime}}(j)$ for $j\in[2L]$. With TCT privacy analysis we only count target hits. The expected number of target hits is $\sum_{j\in[2L]}\min\{\pi_{\varepsilon^{\prime},\delta^{\prime}}(j),1-\pi_{\varepsilon^{\prime},\delta^{\prime}}(j)\}=O(1/\varepsilon^{\prime})$. We obtain a composition over $O(\max\{1/\varepsilon^{\prime},\log(1/\delta)\})$ steps (the $\log(1/\delta)$ is the overhead of TCT). Applying advanced composition and solving for $\varepsilon^{\prime}$ we obtain: $\varepsilon^{\prime}=O((\varepsilon)^{2}\log^{-1}(1/\delta))$ and $\delta^{\prime}=O(\delta)$.
It remains to establish that the claimed property holds. We do so via case analysis. Denote by $C_{t}=\sum_{i=1}^{t}\Delta_{i}$ and $C^{\prime}_{t}=\sum_{i=1}^{t}\Delta^{\prime}_{i}$ the value of the counter $C$ in Algorithm 2 at step $t$ for inputs $\boldsymbol{\Delta}$ and $\boldsymbol{\Delta^{\prime}}$ respectively. Observe that we always have $C_{t}-1\leq C^{\prime}_{t}\leq C_{t}$.
Case $i^{*}\leq K+1$
  • For $t\leq K$, $C^{\prime}_{t}\leq C_{t}\leq K$ and hence $P_{t}=P^{\prime}_{t}=\pi_{(\varepsilon^{\prime},\delta^{\prime})}(0)=0$.
  • For $t=K+1$ (relevant when $i^{*}<K+1$), $C_{t}=K+1$ and hence $P_{t}=\pi_{(\varepsilon^{\prime},\delta^{\prime})}(C_{t}-K)=\pi_{(\varepsilon^{\prime},\delta^{\prime})}(1)=\delta^{\prime}$ and $C^{\prime}_{t}=K$ and hence $P^{\prime}_{t}=\pi_{(\varepsilon^{\prime},\delta^{\prime})}(0)=0$.
  • For $K+1<t\leq K+2L$, $K+2L\geq C_{t}=C^{\prime}_{t}+1\geq K+2$ That is $2\leq C_{t}-K\leq 2L$ and $11\leq C^{\prime}_{t}-K\leq 2L-$ 1 and hence by Claim 3, $Y_{t}\approx_{\varepsilon^{\prime}}Y^{\prime}_{t}$.
  • For $t=K+2L+1$, we have $C_{t}=K+2L+1$ and $C^{\prime}_{t}=K+2L$ and hence $P_{t}=1$ and $P^{\prime}_{t}=1-\delta^{\prime}$.
  • For $t\geq K+2L+1$, $P_{t}=P^{\prime}_{t}=1$.
Case $K+1<i^{*}\leq K+2L+1$
  • For $t<i^{*}$, $P_{t}=P^{\prime}_{t}$.
  • For $i^{*}<t\leq K+2L$, we have $K+2\leq C_{t}=C^{\prime}_{t}+1\leq K+2L$ and $2\leq C_{t}-K\leq K+2L$ and hence by Claim 3, $Y_{t}\approx_{\varepsilon^{\prime}}Y^{\prime}_{t}$.
  • For $t=K+2L+1$ (relevant when $i^{*}<K+2L+1$), we have $C_{t}=K+2L+1$ and hence $P_{t}=1$ and $C^{\prime}_{t}=K+2L$ and hence $P^{\prime}_{t}=1-\delta^{\prime}$.
  • If $t>K+2L+1$, then $C^{\prime}_{t}-K>2L$ and $C_{t}-K>2L+1$ and therefore $P_{t}=P^{\prime}_{t}=1$.
Case $i^{*}>K+2L+1$
In this case for all $t\leq K+2L+1$, $C_{t}=C^{\prime}_{t}$ and hence $P_{t}=P^{\prime}_{t}$ and for all $t>K+2L+1$, $C_{t}-K>2L+1$ and $C^{\prime}_{t}-K>2L$ and hence $P_{t}=P^{\prime}_{t}=1$.
Summarizing all this, in each of the cases we have a composition of at most $22L-$ 2 $\varepsilon^{\prime}$-indistinguishable random variables and at most two $(0,\delta^{\prime})$-indistinguishable random variables.
@@ Appendix C Private Online Prediction for Point Functions
    Instantiate JDP-Mirror (Algorithm 2) with privacy parameters $\varepsilon,\delta$ and with the delay $K=20\cdot k$, where $k=\Theta(\frac{1}{\varepsilon}\log\frac{1}{\delta})$.
@@
        If ${\rm Count}\geq k$ then
@@
          Run an $(\varepsilon,\delta)$-DP algorithm for sparse histograms to identify a point $x^{*}\in\mathcal{X}$ that appears $\Omega(\frac{1}{\varepsilon}\log(1/\delta))$ times in the multiset $\{x_{j}:j\in[t]\text{ and }y_{j}=1\}$. If the algorithm failed to identify such a point then we set $x^{*}\leftarrow\star$ for some special symbol $\star\notin\mathcal{X}$. See, e.g., Bun et al. (2016); Cohen et al. (2021).
@@
[Restatement of Lemma 1.12] Let $\mathcal{X}$ be a domain. The class of point functions over $\mathcal{X}$ admits an $(\varepsilon,\delta)$-DP online predictor with mistake bound $O(\frac{1}{\varepsilon^{2}}\log^{2}\frac{1}{\delta})$.
@@
sketch   The proof is via the construction given in Algorithm 3, which we denote here as algorithm $\mathcal{A}$. The utility analysis of this algorithm is straightforward. We only need to consider the case where the input is consistent with some point function, say with point $x^{*}$. In this case, we argue that the output is identical to that of JDP-Mirror with delay $K$ on inputs $\mathbf{1}(x_{i}=x^{*})$, with $\top$ replaced by $1$ and $\bot$ replaced by 00. Once $\approx\frac{1}{\varepsilon}\log\frac{1}{\delta}$ positive labels have been obtained, we run the private histogram algorithm in order to identify the point $x^{*}$. Since the delay $K$ is much larger than $k$ we have $\text{Flag}=1$. (Note that if JDP-Mirror would have returned a $\top$ before that point, we know that the input is not consistent with a point function.) After JDP-Mirror returns the first $\top$ we get that $\text{Fake}=\textsf{Lap}(\frac{1}{\varepsilon})$ and the algorithm transitions to $\text{Flag}=2$. The output of the algorithm then follows JDP-Mirror and the number of mistakes that is at most $K+O(\varepsilon^{-2}\log^{2}(1/\delta))=O(\varepsilon^{-2}\log^{2}(1/\delta))$.
For the privacy analysis, fix two neighboring input streams that differ on the $i$th labeled example:
$S_{0}$ $=((x_{1},y,1),\dots,(x^{0}_{i},y^{0}_{i}),\dots,(x_{T},y_{T}))$
$S_{1}$ $=((x_{1},y,1),\dots,(x^{1}_{i},y^{1}_{i}),\dots,(x_{T},y_{T})).$
Let $b\in\{0,1\}$ and consider the execution of $\mathcal{A}$ on $S_{b}$. We now claim that the outcome distribution of $\mathcal{A}(S_{b})$, excluding the $i$th output, can be simulated from the outcomes of three private computations on $b$. Specifically, we consider a simulator Sim that interacts with a “helping algorithm”, Helper, who knows the value of the bit $b$ and answers DP queries w.r.t. $b$ as issued by Sim. The simulator Sim first asks Helper to execute the algorithm for histograms on the dataset containing the first $k$ positively labeled points in $S_{b}$ to obtain $x^{*}$. This satisfies $(\varepsilon,\delta)$-DP. The simulator then asks Helper for the outcome vector of executing JDP-Mirror on the stream $\left(I(x_{j},x^{*})\right)_{j=1}^{T}$. The Helper responds with the vector $\vec{b}\in\{\bot,\top,\emptyset\}^{T}$ (excluding the $i$th step). Next, similarly to Step 3(c)ii of the algorithm, let $\hat{L}$ denote the indices of the first $10k$ occurrences of $x^{*}$ in $S_{b}$. The simulator asks Helper for the value of
@@
Now, whenever these two events hold, then the outcome of the simulator is identical to the outcome of $\mathcal{A}(S_{b})$. More specifically, if the algorithm sets ${\rm Flag}=1$ in Step 3iiiC, then the outcome of the simulator is syntactically identical to the output of the algorithm. Otherwise (if we set ${\rm Flag}=0$ in Step 3iiiC), then there must be at least $9k$ occurrences of $(x^{*},0)$ in $S_{b}$ before the $k$th occurrence of $(x^{*},1)$. Hence, in this case we have that $\widehat{{\rm Fake}}\geq\frac{1}{\varepsilon}\log\frac{1}{\delta}$. So both the algorithm and the simulator output the all zero vector.